# Predictable Scale: Part I — Optimal Hyperparameter Scaling Law in Large Language Model Pretraining

### v4

@@ 4.2.2 Scaling Laws
Building upon the insights gained from Section 4.2.1, we delve into the scaling behavior of optimal hyperparameters. Specifically, we investigate how the optimal LR scales with $N$ and $D$, and how the optimal BS scales with $D$. Our empirical observations, particularly when visualized on a log-log scale, reveal a strong linear trend, suggesting a power-law relationship. Based on this, the scaling law for hyperparameters can be described by the following power-law relationships:
@@
where the parameters $c,\alpha,\beta,d$, and $\gamma$ are five constants, the values of which will be determined through fitting in Section 4.2.3. It is particularly noteworthy that our proposed scaling law demonstrates significant generality, meaning it is applicable across diverse architectures $\mathbb{A}$ and data distributions $\mathbb{D}$. This aspect of generality will be further elaborated upon in Section 5.
@@
(a) Scaling laws with $D$ for different $N$
(b) Scaling laws with $N$ for different $D$
Figure 6: (a) Scatter points indicate empirical optimal learning rate vs. batch size for model scale $N$; (b) Analogous results for dataset scale D. Curves show our hp-scaling law predictions, with shaded regions representing parameter uncertainty bounds from the sampling-based fitting strategy. Both plots use double logarithmic scaling (1912 training samples).
4.2.3 Fitting Methods
Building upon the HP scaling law from Section 4.2.1, we transform the power laws in Eq. (6) into the linear form:
@@ 4.2.3 Fitting Methods
In this way, we can employ Ordinary Least Squares to fit the unknown parameters $\log c,\alpha,\beta,\log d$ and $\gamma$. Specifically, we set up 7 groups of experiments with different $N$ and $D$ as shown in Appendix A.2. Following Hoffmann et al. (2022), we fit the optimal LR and BS with the experimentally predicted LR and BS. We averaged the results of these 1000 bootstrap samples to obtain the intermediate final parameters. This averaged result is what we present in Tab. 2. Furthermore, the variability across these 1000 bootstrap samples is depicted as the shaded regions in Fig. 6, providing an indication of the uncertainty associated with the fitted results. These shaded regions allow us to visually assess the robustness and confidence of the optimal LR and BS values derived from our procedure.
4.2.4 Experimental Comparison with Existing Approaches
Since we have obtained the fitted scaling laws, we directly extrapolate them to the test point ($N=1$B and $D=100$B) for comparison with other methods, noting that these are out-of-sample extrapolations beyond the fitting range. As shown in Fig. 1 and Fig. 11, our method predicts solutions closest to the global minimum, while other approaches (indicated by dashed lines) remain distant from the global minimum. This discrepancy arises because previous methods only considered LR fitting without jointly modeling both LR and BS. The approach in DeepSeek-AI et al. (2024a) enforces a fixed compute budget assumption, which constrains the feasible $(N,D)$ sampling range and consequently reduces fitting accuracy. Additionally, as discussed in Section 4.1.3, existing methods typically relate final LR to initial LR through a fixed multiplier. This assumption leads to excessively large final LRs when initial LRs are large, ultimately impairing convergence. While Porian et al. (2024) achieves comparable results to ours through similar $\mathrm{min\_lr}$ constraints, their method exhibits instability due to incomplete consideration of hyperparameter interactions with model dimension $D$, particularly at the subsequent MoE experiments  5.2 and data recipe experiments 5.3, where the $D/N$ ratios are relatively small, their methods predict learning rates and batch sizes that fall completely outside reasonable ranges, leading to training instability.
Key Takeaways HP Scaling Law: We observed that the optimal LR follows a power law with respect to both $N$ and $D$, while the optimal BS follows a power law only with respect to $D$ and remains relatively invariant to $N$. Based on these observations, we derived our formula. Experimental Comparison: Relative to other approaches, our approach involved a substantial resource allocation to comprehensively analyze the dependence of LR and BS on $N$ and $D$. This resulted in a marked improvement in performance.
5 Universal HP Scaling laws: Empirical Validation Across Architectural Heterogeneity and Data Recipes
5.1 Topological Invariance Across Varied Model Shape
Figure 7: Topological Invariance Across Varied Model Shape. $d_{\text{model }}$, $d_{\mathrm{ff}}$, $N_{\text{layer}}$, and $N_{\text{head}}$ denote the hidden dimension, feed-forward network hidden size, number of attention heads, and number of Transformer layers, respectively.
@@ 5.1 Topological Invariance Across Varied Model Shape
For each of these 6 model shapes, we perform extensive hyperparameter tuning to identify the optimal LR and BS. The results reveal a striking pattern: the optimal LR and BS values for all configurations (highlighted within the solid-line box) consistently fall within a well-defined and narrow range (enclosed by the dashed-line box). This consistency holds across all model shape combinations, despite significant variations in architectural topology. These empirical findings provide strong evidence supporting our hypothesis that the HP scaling law exhibits statistical invariance with respect to changes in model topology. In other words, while the architectural components—including depth (number of layers), attention mechanism complexity (number of attention heads), and feedforward network width—may vary, the fundamental scaling relationships governing LR, BS, model size $N$, and dataset size $D$ remain unchanged.
5.2 Sparsity-Independent in MoE
The HP scaling law has been extensively studied for dense Transformers, but its applicability to sparse architectures remains uncertain. Mixture-of-Experts (MoE) Shazeer et al. (2017); Fedus et al. (2022) is a widely used sparse model that activates only a subset of parameters per token, introducing fundamental structural differences from dense models. This raises the question of whether the HP scaling law can be generalized to MoE settings. To investigate this, we conducted experiments on MoE models across 16 different sparsity levels and model shapes (refer to Tab. 5 in the appendix A.2). These settings allow us to examine how the scaling law behaves under different levels of sparsity. We evaluate multiple existing scaling methods under this framework.
As shown in Fig. 8, our approach consistently achieves a relative prediction error within 0.5% across all sparsity levels, significantly outperforming competing methods. In contrast, the DeepSeek Formula yields a relative error over four times larger, indicating its reduced accuracy in MoE settings. While Eq.(1) achieves comparable accuracy in LR prediction, it fails to predict BS. In contrast, our method provides a more comprehensive framework, successfully predicting multiple hyperparameters. These results demonstrate that the HP scaling law extends beyond dense architectures and remains effective for sparse models like MoE, regardless of sparsity level. This suggests that the underlying principles of scaling laws emerge from broader optimization and capacity constraints rather than being specific to dense parameterization. Our findings reinforce the general applicability of HP scaling laws and their potential to guide efficient scaling in diverse neural architectures $\mathbb{A}$.
@@ 5.2 Sparsity-Independent in MoE
5.3 Data-Distribution Robustness
@@
To rigorously assess the robustness of our HP scaling law across varied data distributions $\mathbb{D}$, we design three distinct data distributions, progressively diverging from the original composition, as detailed in Appendix Tab. 6:
  1. 1.
  2. 2.
  3. 3.
As shown in Fig. 9, our formula maintains remarkable predictive accuracy across all three distributions, with relative prediction errors within 0.25% of the global minimum. This performance consistently surpasses alternative approaches, which exhibit larger deviations. These results highlight two crucial insights:
  • The HP scaling law demonstrates statistical invariance across linguistic and structural changes in the dataset, supporting its generalizability beyond standard natural language distributions.
  • The predicted optimal hyperparameters remain stable even with highly heterogeneous training data, reinforcing our approach’s robustness.
These findings are particularly significant for designing scalable and adaptable training paradigms applicable across diverse deployment scenarios with varying dataset characteristics.
Key Takeaways Topological Invariance: Our HP scaling laws exhibit statistical invariance in the scaling constants for LR and BS with respect to model scale $N$ and data size $D$, even when varying topological features of model architectures. Sparsity Independence: The HP scaling law extends beyond dense Transformers and remains effective for sparse MoE models, with our approach achieving superior prediction accuracy across different sparsity levels, reinforcing the broader applicability of scaling laws in diverse neural architectures $\mathbb{A}$. Data-Distribution Robustness: The HP scaling law shows robustness across diverse data distributions $\mathbb{D}$.
@@ 6 Conclusions
In this paper, we provide a crucial advancement in efficient hyperparameter optimization for LLMs. By empirically unveiling and rigorously validating universal scaling laws for learning rate and batch size—underpinned by the discovery of loss landscape convexity—we move beyond computationally expensive grid searches and limited transfer methods. Our robust HP scaling laws, supported by an unprecedentedly large empirical study and open-sourced resources, empower the community with a practical and generalizable approach for navigating the hyperparameter configuration space in LLM pretraining, thereby facilitating more efficient and scalable LLM development.
Limitations
Acknowledgments
The work was supported by National Science and Technology Major Project of China (2023ZD0121300).
@@ Appendix A Appendix
A.1 Model Scale Dominates Optimal Hyperparameter Selection Over Computational Complexity
$d_{{model}}$ $d_{{ff}}$ $N_{{head}}$ $N_{{layer}}$ $lr$ $bs$ $D$ $N$ $M$
@@ A.1 Model Scale Dominates Optimal Hyperparameter Selection Over Computational Complexity
1280 12264 10 8 1.95E-03 262,144 8.00E+09 4.29E+08 2.83E+09
1280 6280 10 14 1.38E-03 262,144 8.00E+09 4.29E+08 3.02E+09
1536 9600 12 8 9.77E-04 131,072 8.00E+09 4.29E+08 2.88E+09
1536 7264 12 10 1.38E-03 262,144 8.00E+09 4.29E+08 2.95E+09
1536 4608 12 14 9.77E-04 131,072 8.00E+09 4.29E+08 3.10E+09
2048 6000 16 8 9.77E-04 262,144 8.00E+09 4.29E+08 2.98E+09
2048 4256 16 10 9.77E-04 262,144 8.00E+09 4.29E+08 3.08E+09
2048 2256 16 14 9.77E-04 262,144 8.00E+09 4.29E+08 3.28E+09
@@
1280 12608 10 8 1.38E-03 262,144 8.00E+09 4.40E+08 2.89E+09
1280 5888 10 14 1.38E-03 262,144 8.00E+09 4.08E+08 2.89E+09
1536 9656 12 8 1.38E-03 262,144 8.00E+09 4.31E+08 2.89E+09
1536 7040 12 10 1.38E-03 262,144 8.00E+09 4.19E+08 2.89E+09
1536 4056 12 14 9.77E-04 262,144 8.00E+09 3.94E+08 2.89E+09
2048 5704 16 8 9.77E-04 262,144 8.00E+09 4.15E+08 2.89E+09
2048 3744 16 10 6.91E-04 131,072 8.00E+09 3.98E+08 2.89E+09
2048 1504 16 14 6.91E-04 131,072 8.00E+09 3.64E+08 2.89E+09
Table 3: Model configurations for constant $N$ and constant $M$ experiments. The first group (top) maintains constant parameter count $N$approx 4.29E+08, while the second group (bottom) maintains constant computational complexity $M$approx 2.89E+09. $M$: non-embedding FLOPs/token.
@@
Tab. 3 presents the detailed configurations and results for both experimental groups. For each model, we systematically varied the hidden dimension ($d_{{model}}$), feed-forward dimension ($d_{{ff}}$), number of attention heads ($N_{{head}}$), and number of layers ($N_{{layer}}$) while maintaining either constant $N$ or $M$. The embedding dimension ($D$) was fixed at 8.00E+09 across all experiments.
To visualize the impact of hyperparameters across different architectural configurations, we generated heatmaps of the loss landscape with respect to LR and BS in Fig. 7 and 10. The heatmaps reveal consistent patterns in the optimal hyperparameter regions across different architectural configurations within each experimental group.
@@
The experimental results reveal several key findings: (i) Models with constant $N$ demonstrate remarkably consistent optimal hyperparameter regions, with minimal variation in minimum loss values (ranging from 2.4294 to 2.4776) despite significant architectural differences. (ii) The constant M experiments show slightly more variation in optimal hyperparameter regions and minimum loss values (ranging from 2.4346 to 2.5089), suggesting that parameter count $N$ may be a more robust indicator for hyperparameter selection than computational complexity $M$. (iii) Across both experimental groups, the optimal learning rates typically fall within a narrow range (6.91E-04 to 1.95E-03), and batch sizes cluster around either 131,072 or 262,144, regardless of the specific architectural configuration.
These findings strongly suggest that the fundamental scale metrics, particularly the parameter count $N$, are more influential in determining optimal hyperparameter settings than specific architectural choices. This observation motivates our discussion of hyperparameter scaling laws in relation to $N$ in Section 4.2.
A.2 Model Structural Parameters And Results

### v6

@@ 3.4.1 Scaling Laws
In accordance with Definition 1, we experimentally derive the LR and BS by keeping other variables fixed. This section focuses on elucidating the relationships between these empirically determined hyperparameters and $N$ and $D$. For the parameter count $N$, we set up seven experiments spanning 60M, 120M, 210M, 270M, 430M, 540M, and 1B parameters. As demonstrated in Fig. 5(a), our experiments reveal a positive correlation between optimal LR and BS and the data scale $D$ for each value of $N$. Furthermore, we conducted experiments across five different data scales $D$: 2B, 4B, 8B, 20B, and 100B tokens. Notably, we specifically reserved the 1B parameter and 100B token settings as test points to validate our findings, as discussed in Sec. 3.4.3. As visualized in Fig. 5(b), we find that for each data scale $D$, the optimal LR increases with model size $N$. Notably, our findings indicate that optimal BS is largely independent of $N$.
Building upon the above insights, we delve into the scaling behavior of optimal hyperparameters. Specifically, we investigate how the optimal LR scales with $N$ and $D$, and how the optimal BS scales with $D$. Our empirical observations, particularly when visualized on a log-log scale, reveal a strong linear trend, suggesting a power-law relationship. Based on this, the scaling law for hyperparameters can be described by the following power-law relationships:
@@
where $c,\alpha,\beta,d$, and $\gamma$ are constants, the values of which will be determined through fitting in Sec. 3.4.2. Notably, the proposed form $B(D)$ assumes that the optimal batch size is independent of $N$. This assumption is statistically validated through regression analysis in Appendix A.5. It is particularly noteworthy that our proposed scaling law demonstrates significant generality, meaning it is applicable across diverse architectures $\mathbb{A}$ and data recipes $\mathbb{D}$. This aspect of generality will be further elaborated upon in Sec. 3.5.
(a) Scaling laws with dataset size for different number of parameters
(b) Scaling laws with number of parameters for different dataset size
Figure 5: Empirical vs. predicted HP scaling. (a) Scatter points indicate empirical optimal learning rate vs. batch size for model scale $N$; (b) Analogous results for dataset scale D. Curves show our hp-scaling law predictions, with shaded regions representing parameter uncertainty bounds from the sampling-based fitting strategy. Both plots use double logarithmic scaling (1,912 training samples).
3.4.2 Fitting Methodology
@@
Building on Step Law from Sec. 3.4.1, we transform the power laws in Eq. (6) into the linear form:
@@
In this way, we can employ Ordinary Least Squares to fit the unknown parameters $\log c,\alpha,\beta,\log d$ and $\gamma$. Specifically, we set up 7 groups of experiments with different $N$ and $D$ as shown in Appendix A.3. Following hoffmann_training_2022 , we fit the optimal LR and BS with the experimentally predicted LR and BS. We averaged the results of these 1,000 bootstrap samples to obtain the intermediate final parameters. This averaged result is what we present in Tab. 2. Furthermore, the variability across these 1,000 bootstrap samples is depicted as the shaded regions in Fig. 5, providing an indication of the uncertainty associated with the fitted results. These shaded regions allow us to visually assess the robustness and confidence of the optimal LR and BS derived from our procedure.
3.4.3 Comparisons with Existing Approaches
With the fitted scaling laws, we directly extrapolate to the test point ($N=1$B, $D=100$B) for comparison. As shown in Fig. 1 and detailed in Appendix A.7, our method consistently finds solutions near the global optimum. In contrast, other approaches show significant deviation.This gap mainly comes from their modeling limitations.
Previous methods typically fit learning rate alone, without jointly modeling batch size. For example, DeepSeek LawDeepSeek-AI2024 assumes a fixed compute budget, which restricts the $(N,D)$ range and degrades fitting accuracy. As discussed in Sec. 3.3.2, many approaches fix the final LR as a constant multiple of the initial LR. This often results in overly large final LRs for large initial values, leading to poor convergence. Although Porian Law Porian2024 mitigate this with a minimum LR constraint, their method lacks stability. It does not fully account for interactions between hyperparameters and model dimension $D$. This becomes more problematic in MoE (Sec. 3.5.2) and data recipe (Sec. 3.5.3), where small $D/N$ ratios cause their predicted LR and BS to fall outside stable ranges.
We further test on a 6.5B-parameter MoE model with two dataset sizes: $D=1.0\times 10^{10}$ and $1.3\times 10^{11}$ (see configurations 17 and 18 in Appendix A.3). For each setting, we use Porian Law, DeepSeek Law, and our Step Law to predict LR and BS from $(N,D)$, and train accordingly. We exclude $\mu$Yang2022 as it only predicts LR and does not handle BS.Table 3 shows that our Step Law consistently achieves the lowest validation smooth loss. This confirms that our method generalizes well to large, sparse MoE architectures and remains robust under diverse training regimes.
Table 3: Comparison of predicted HPs and validation smooth loss on a 6.5B MoE model.
Method Learning Rate (LR) Token-wise Batch Size (BS) Smooth Loss
Dataset size $D=1.0\times 10^{10}$
Porian Law Porian2024
$1.08\times 10^{-3}$ 6029312 2.4352
DeepSeek Law DeepSeek-AI2024
$1.06\times 10^{-3}$ 851968 2.2891
Step Law (Ours) $2.12\times 10^{-4}$ 294912 2.2700
Dataset size $D=1.3\times 10^{11}$
Porian Law Porian2024
$1.08\times 10^{-3}$ 6029312 1.9772
DeepSeek Law DeepSeek-AI2024
$7.70\times 10^{-4}$ 1966080 1.9511
Step Law (Ours) $4.74\times 10^{-4}$ 1310720 1.9479
3.5 Validation Cross Architecture and Data
3.5.1 Topological Invariance Across Varied Model Shapes
Figure 6: Consistent Hyperparameter Optima Across Diverse Model Topologies. Smooth loss landscapes across learning rate (LR) and batch size (BS) for six Transformer architectures, each with approximately 430M parameters but differing in shape configurations, varying hidden dimension ($d_{\text{model}}$), feedforward size ($d_{\mathrm{ff}}$), number of layers ($N_{\text{layer}}$), and attention heads ($N_{\text{head}}$). Despite these architectural differences, the optimal LR-BS region (solid box) consistently resides within a narrow band (dashed box), indicating topological invariance in hyperparameter scaling behavior.
@@ 3.5.1 Topological Invariance Across Varied Model Shapes
For each of the 6 model shapes, we conduct extensive hyperparameter tuning to identify the optimal LR and BS. The results show a clear pattern: across all configurations, the optimal LR and BS consistently fall within a narrow range, despite large variations in model topology. This consistency supports our hypothesis that Step Law is invariant to architectural changes. In particular, variations in depth, attention heads, or feedforward width do not alter the scaling relationships among LR, BS, model size $N$, and dataset size $D$.
3.5.2 Sparsity-Robustness in MoE Models
Step Law has been extensively studied for dense model, but its applicability to sparse architectures remains uncertain. MoE shazeer2017outrageously ; fedus2022switch activates only a subset of parameters per token, making it structurally different from dense models. This raises the question of whether Step Law can be generalized to MoE settings. To investigate this, we conducted experiments on MoE models across 16 different sparsity levels and model shapes (refer to Tab. 6 in the Appendix A.3). These settings allow us to examine how the scaling law behaves under different levels of sparsity. We evaluate multiple existing scaling methods under this framework.
As shown in Fig. 7, our approach consistently achieves a relative prediction error within 0.5% across all sparsity levels, significantly outperforming competing methods. In contrast, the DeepSeek Formula yields a relative error over four times larger, indicating its reduced accuracy in MoE settings. In contrast, our method provides a more comprehensive framework, successfully predicting multiple hyperparameters. Additional MoE experiment details and full results across 16 sparsity levels are provided in  A.8. These results demonstrate that the Step Law extends beyond dense architectures and remains effective for MoE, regardless of sparsity level. This suggests that the underlying principles of scaling laws emerge from broader optimization and capacity constraints rather than being specific to dense parameterization. Our findings reinforce the general applicability of Step Laws and their potential to guide efficient scaling in diverse neural architectures $\mathbb{A}$.
@@ 3.5.2 Sparsity-Robustness in MoE Models
3.5.3 Data Distribution Generalization
@@
To rigorously assess the robustness of our Step Law across varied data recipes $\mathbb{D}$, we design three distinct data recipes, progressively diverging from the original composition, as detailed in Appendix Tab. A.4: adding Chinese (Bilingual Corpus), adding code (Code Integration), and mostly code (Code-Dominant).As shown in Fig. 8, our formula achieves high predictive accuracy across all three distributions, with relative errors within 0.25% of the global minimum. This consistently outperforms alternative methods, which show larger deviations.
These results reveal two key findings: (a) The Step Law exhibits statistical invariance under both linguistic and structural variations, suggesting strong generalizability beyond standard language distributions; (b) The predicted optimal hyperparameters remain stable across highly heterogeneous training data, demonstrating the robustness of our approach. These findings are particularly significant for designing scalable and adaptable training paradigms applicable across diverse deployment scenarios with varying dataset characteristics.
@@ 4 Conclusions
In this paper, we provide a crucial advancement in efficient hyperparameter optimization for LLMs. By empirically unveiling and rigorously validating universal scaling laws for learning rate and batch size—underpinned by the discovery of loss landscape convexity—we move beyond computationally expensive grid searches and limited transfer methods. Our robust HP scaling laws, supported by an unprecedentedly large empirical study and open-sourced resources, empower the community with a practical and generalizable approach for navigating the hyperparameter configuration space in LLM pre-training, thereby facilitating more efficient and scalable LLM development.
@@ Appendix A Appendix
A.1 Notation
We use the following notation:
  • $\mathcal{L}$: Cross-entropy loss.
  • $D$: Dataset size in tokens.
  • $N$: Number of non-embedding parameters in the model.
  • $\hat{N}$: Total number of parameters in the model, which excludes embedding layer but includes the models head layer.
  • $C$: Compute budget in FLOPs.
  • $N_{\text{layer}}$: Number of layers in the Transformer model.
  • $d_{\text{ff}}$: Dimension of the feed-forward network hidden layer in the Transformer.
  • $d_{\text{model}}$: Hidden dimension of the Transformer model.
  • $N_{\text{head}}$: Number of attention heads in the Transformer model.
  • LR: Learning rate.
  • BS: Batch size (in tokens).
  • $B(N,D)$: Optimal batch size (in tokens) for given parameter $N$ and dataset size $D$.
  • $\mathbb{A}$: Model architecture space defined by $N_{\text{layer}}$, $d_{\text{ff}}$, $d_{\text{model}}$, and $N_{\text{head}}$.
  • $\mathbb{D}$: Training data distribution governed by the data-generating probability distribution.
A.2 Model Scale Dominates Optimal Hyperparameter Selection Over Computational Complexity
$d_{{model}}$ $d_{{ff}}$ $N_{{head}}$ $N_{{layer}}$ $\eta(N,D)$ $B(N,D)$ $D$ $N$ $M$
@@ A.2 Model Scale Dominates Optimal Hyperparameter Selection Over Computational Complexity
1280 12264 10 8 $1.95\times 10^{-3}$ $262{,}144$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $2.83\times 10^{9}$
1280 6280 10 14 $1.38\times 10^{-3}$ $262{,}144$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $3.02\times 10^{9}$
1536 9600 12 8 $9.77\times 10^{-4}$ $131{,}072$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $2.88\times 10^{9}$
1536 7264 12 10 $1.38\times 10^{-3}$ $262{,}144$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $2.95\times 10^{9}$
1536 4608 12 14 $9.77\times 10^{-4}$ $131{,}072$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $3.10\times 10^{9}$
2048 6000 16 8 $9.77\times 10^{-4}$ $262{,}144$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $2.98\times 10^{9}$
2048 4256 16 10 $9.77\times 10^{-4}$ $262{,}144$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $3.08\times 10^{9}$
2048 2256 16 14 $9.77\times 10^{-4}$ $262{,}144$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $3.28\times 10^{9}$
@@
1280 12608 10 8 $1.38\times 10^{-3}$ $262{,}144$ $8.00\times 10^{9}$ $4.40\times 10^{8}$ $2.89\times 10^{9}$
1280 5888 10 14 $1.38\times 10^{-3}$ $262{,}144$ $8.00\times 10^{9}$ $4.08\times 10^{8}$ $2.89\times 10^{9}$
1536 9656 12 8 $1.38\times 10^{-3}$ $262{,}144$ $8.00\times 10^{9}$ $4.31\times 10^{8}$ $2.89\times 10^{9}$
1536 7040 12 10 $1.38\times 10^{-3}$ $262{,}144$ $8.00\times 10^{9}$ $4.19\times 10^{8}$ $2.89\times 10^{9}$
1536 4056 12 14 $9.77\times 10^{-4}$ $262{,}144$ $8.00\times 10^{9}$ $3.94\times 10^{8}$ $2.89\times 10^{9}$
2048 5704 16 8 $9.77\times 10^{-4}$ $262{,}144$ $8.00\times 10^{9}$ $4.15\times 10^{8}$ $2.89\times 10^{9}$
2048 3744 16 10 $6.91\times 10^{-4}$ $131{,}072$ $8.00\times 10^{9}$ $3.98\times 10^{8}$ $2.89\times 10^{9}$
2048 1504 16 14 $6.91\times 10^{-4}$ $131{,}072$ $8.00\times 10^{9}$ $3.64\times 10^{8}$ $2.89\times 10^{9}$
Table 4: Model configurations and results for constant $N$ and constant $M$ experiments. The first group (top) maintains constant parameter count $N$approx $4.29\times 10^{8}$, while the second group (bottom) maintains constant computational complexity $M$approx $2.89\times 10^{9}$. $M$: non-embedding FLOPs/token.
@@
Tab. 4 presents the detailed configurations and results for both experimental groups. For each model, we systematically varied the hidden dimension ($d_{{model}}$), feed-forward dimension ($d_{{ff}}$), number of attention heads ($N_{{head}}$), and number of layers ($N_{{layer}}$) while maintaining either constant $N$ or $M$. The embedding dimension ($D$) was fixed at $8.00\times 10^{9}$ across all experiments.
To visualize the impact of hyperparameters across different architectural configurations, we generated heatmaps of the loss landscape with respect to LR and BS in Fig. 6 and 9. The heatmaps reveal consistent patterns in the optimal hyperparameter regions across different architectural configurations within each experimental group.
@@
[…]
These findings strongly suggest that the fundamental scale metrics, particularly the parameter count $N$, are more influential in determining optimal hyperparameter settings than specific architectural choices. This observation motivates our discussion of hyperparameter scaling laws in relation to $N$ in Sec. 3.4.
A.3 Model Structural Parameters
Table 5 and Table 6 summarizes the precise architectural settings for all models evaluated in our study. In the dense model group (Models 1-18), we cover models ranging from $2.15\times 10^{8}$ to $1.07\times 10^{9}$ total parameters by varying the hidden dimension ($d_{\mathrm{model}}$), feed-forward width ($d_{\mathrm{ff}}$), number of attention heads ($N_{\mathrm{head}}$) and number of layers ($N_{\mathrm{layer}}$). Each entry also lists the corresponding dataset size $D$ used during training. In the MoE model group (Models 1-16), we hold the overall parameter count fixed at approximately $2.15\times 10^{9}$ while sweeping dataset size $D$ from $2\times 10^{9}$ up to $2\times 10^{10}$. We further vary the number of experts ($N_{\mathrm{expert}}$), per-expert hidden size ($d_{\mathrm{moe}}$), top-$k$ routing, and the resulting active parameter count ($N_{a}$). This systematic variation allows direct comparison of dense versus sparse Mixture-of-Experts architectures under matched compute budgets and data scales.